# Channel Measurement, Modeling, and Simulation for 6G: A Survey and Tutorial

### v3

@@ Abstract
The sixth generation (6G) mobile communications have attracted substantial attention in the global research community of information and communication technologies (ICT). 6G systems are expected to support not only extended 5G usage scenarios, but also new usage scenarios, such as integrated sensing and communication (ISAC), integrated artificial intelligence (AI) and communication, and communication and ubiquitous connectivity. To realize this goal, channel characteristics must be comprehensively studied and properly exploited, so as to promote the design, standardization, and optimization of 6G systems. In this paper, we first summarize the requirements and challenges in 6G channel research. Our focus is on channels for five promising technologies enabling 6G, including terahertz (THz), extreme MIMO (E-MIMO), ISAC, reconfigurable intelligent surface (RIS), and space-air-ground integrated network (SAGIN). Then, a survey of the progress of the 6G channel research regarding the above five promising technologies is presented in terms of the latest measurement campaigns, new characteristics, modeling methods, and research prospects. Moreover, a tutorial on the 6G channel simulations is presented. We introduce the BUPTCMCCCMG-IMT2030, a 6G link-level channel simulator, developed based on the ITU/3GPP 3D geometry-based stochastic model (GBSM) methodology. The simulator supports the channel simulation of the aforementioned 6G potential technologies. To facilitate the use of the simulator, the tutorial encompasses the design framework, user guidelines, and application examples. This paper offers in-depth, hands-on insights into the best practices of channel measurements, modeling, and simulations for the evaluation of 6G technologies, the development of 6G standards, and the implementation and optimization of 6G systems.
@@
6G, channel modeling, channel measurement, channel simulation, terahertz, MIMO, integrated sensing and communication, reconfigurable intelligent surface, space-air-ground integrated network
@@ I-A Overview of 6G
As the fifth generation (5G) mobile communication networks have been deployed worldwide, studies on the sixth generation (6G) mobile communication technologies, which are expected to be rolled out for 2030 and beyond, are being vibrantly conducted. Several 6G research projects and organizations have been established. Representative examples include the Finnish “6Genesis Flagship Program” [1], the European “Hexa-X” project [2], the Chinese “IMT-2030 (6G) Promotion Group” [3], and the North American “Next G Alliance” [4]. Simultaneously, international standardization organizations have launched several key initiatives. For instance, in 2022, the International Telecommunication Union (ITU) identified future technology trends [5] for International Mobile Telecommunications (IMT) systems toward 2030 and beyond. Very recently, the recommendation framework for IMT-2030, which is commonly known as the 6G vision, was finalized [6]. This means that the global 6G standardization campaign has been formally initiated. With the recommendation framework, ITU establishes the overall objectives, procedures and timelines for the development of 6G mobile communication standards. Following the ITU requirements and guidelines, the third Generation Partnership Project (3GPP) is expected to start the drafting of the 6G specifications in approximately 2025, to prepare a technology proposal for the ITU radiocommunication sector (ITU-R) afterwards, and to support the completion of 6G standardization no later than 2030.
@@
6G aims to establish a more inclusive information society than 5G. Its goal is to intelligently connect humans, machines, and various other things [6]. Some key drivers of 6G have been discussed in [5]. These drivers can be understood from the perspective associated with continuously growing mobile traffic and emerging applications. Specifically, it is forecast that by 2030, global mobile traffic will reach 5016 exabytes (EB) per month [7], which means a growth of approximately 80 times compared with that of 2020. The potential emerging applications of 6G [5] can be categorized into several main usage scenarios [6]. One extends enhanced mobile broadband (eMBB), which was introduced in 5G, to target immersive applications, such as holographic communications, tactile and haptic internet applications, and extended reality (XR). Similarly, hyperreliable and low-latency communications and massive communications are extensions of 5G usage scenarios of ultrareliable and low latency communications (URLLC) and massive machine type communications (mMTC), respectively. In addition, three other usage scenarios have been identified, i.e., ubiquitous connectivity, integrated artificial intelligence (AI) and communication, and integrated sensing and communication (ISAC) [6]. The applications of integrated AI/sensing and communication include perceptive mobile networks, digital twins, and the proliferation of networked AI. As a result, the capabilities of 6G are enhanced across multiple dimensions, including data rate, spectrum efficiency, reliability, latency, mobility, and connection density [8, 9, 10, 11, 12, 13, 14, 6]. Moreover, new capabilities are considered, such as all-domain coverage, sensing-related and AI-related capabilities, sustainability and computing power [6].
@@ I-B Requirements and Challenges of 6G Channel Research
The radio channel channel is the transmission medium between a transmitter (Tx) and receiver (Rx). It plays a crucial role in determining the ultimate performance limit of mobile communications systems [15, 16, 17]. As shown in Fig. 1, 6G communication systems will incorporate various novel technological components [5, 16, 18, 10, 19, 20, 21, 8, 22], such as the extension of the frequency band to terahertz (THz) [23, 24, 25, 26, 17, 27], the extension of the antenna array to the extreme MIMO (E-MIMO) [9, 22, 28, 29], the integration of sensing into the communication system (i.e. ISAC) [30, 31, 32], the introduction of a programmable wireless environment by using reconfigurable intelligent surface (RIS) [33, 34, 35], and the extension of coverage to space-air-ground integrated network (SAGIN) [36, 37, 38]. Each of these may result in distinct requirements and characteristics for the wireless channel, as summarized below.
@@
I-B1 Higher frequency band, propagation attenuation, and channel sparsity
The deployment of 5G has already achieved highly efficient utilization of the sub-6 GHz frequency band and harnessed millimeter wave (mmWave) frequencies to deliver mobile devices with unparalleled multi-Gbps data rates [39]. The THz frequency band (0.3-3 THz according to ITU radio regulations or 0.1-10 THz as a commonly used wider range) possesses an atmospheric transmission window with low molecular absorption losses, and abundant bandwidth having been allocated to terrestrial mobile communication services at the World Radiocommunication Conference (WRC) in 2019 [40]. As the era of 6G begins, the demand for a peak data rate of 1 Tbps requires a comprehensive exploitation of the THz frequency band [16]. Prototype demonstration and testing systems for THz communication have already been developed to achieve high-speed transmission. For instance, high-capacity wireless transmission at 1.44 Tbps has been achieved within the THz frequency band of 135-170 GHz [41]. However, it is essential to recognize that the THz band has distinctive characteristics in contrast to the sub-6 GHz and mmWave frequency bands. The shorter wavelength ($\leq$ 3 mm) signals in the THz band experience the roughness of the objects in the environment. Previously smooth surfaces, such as walls and tree leaves, appear rougher at THz band. The specular reflection may lose its privileged position, allowing diffuse scattering to become the predominant reflection behavior. In general, the THz wave experiences severe transmission losses and molecular absorption, thereby causing changes in the path loss characteristics and the emergence of sparsity. Sparsity refers to the condition where only a small number of multipath components (MPCs) carry the majority of the received signal power. Currently, the ITU/3GPP channel model [42, 43] only supports frequencies up to 100 GHz for 5G. Therefore, it is urgent to establish a comprehensive THz channel model, which is capable of covering the typical scenarios (such as indoor hotspot (InH) and urban microcell (UMi)) and incorporating the distinctive THz channel characteristics (such as high path loss and sparsity).
I-B2 Larger-scale antenna array, near-field effect, and spatial non-stationarity
It is the 1970’s that researchers suggested to introduce MIMO technologies into wireless communication systems. In 1995, Telatar theoretically derived the MIMO channel capacity under the fading condition [44]. It was demonstrated that the MIMO technology can substantially increase the capacity of wireless communication system. Then, the MIMO technology was introduced into the fourth generation (4G) mobile communication system, which can support eight array elements and highly improve the system spectrum efficiency [45]. However, the channel model used in 4G systems only supports the azimuth dimension. With the proposal of massive MIMO, three-dimensional (3D) MIMO [46] is a typical massive antenna realization. Therefore, the elevation dimension was introduced into the ITU/3GPP 5G standard channel model[42, 43]. The 5G channel model captures the elevation-azimuth-time-frequency channel characteristics, and meets the requirements of 3D MIMO and massive MIMO technologies. In the 6G era, the number of array elements is expected to increase significantly, reaching a range from a few hundreds to thousands. As a result, the E-MIMO technology based on the much larger-scale antenna arrays is envisaged to be deployed. The E-MIMO with 1024$\times$1024 array elements can offer abundant spatial degrees of freedom for supporting 100 times peak data rate improvement in 6G mobile communication systems [47]. However, some new channel characteristics emerge in E-MIMO systems. To elaborate a little further, when the propagation distance is shorter than the far-field Rayleigh distance, the plane wave assumption is invalid in the near-field region [48, 49]. Therefore, the the near-field spherical wave and spatial non-stationary (SnS) characteristics [26, 50] cannot be ignored in E-MIMO communication systems. As shown in Fig. 1, the E-MIMO can be modeled using near-field spherical propagation in the near-field region. In addition, with the expansion of physical array aperture, different array elements see different scatterers, resulting in SnS on the array. However, these characteristics are not covered by the ITU/3GPP 5G channel model framework. Therefore, it is necessary to investigate the impact of the near-field spherical wave and the SnS characteristic through channel measurements and simulations for 6G channel modeling.
I-B3 ISAC, simultaneous presence of sensing channel and communication channel
@@
ISAC is one of the candidate technologies for 6G[51]. It is capable of exploiting the potential sensing ability of base stations (BSs) to obtain environmental information to assist communication, thus benefiting beam alignment, directional coverage, etc. At the same time, multipath information in communication signals can also enhance sensing ability. In ISAC systems, sensing aims to obtain the state information, such as the speed and location, of the targets in the particular environment through echoes and the collaboration between BSs [52]. Therefore, in the study of ISAC channels, it is essential to introduce the radar cross section (RCS), a target-related feature, and consider the impact of the environment on the sensing channel [53]. Moreover, due to the simultaneous use of the propagation environment by the communication and sensing signals, some shared propagation scatterers may be experienced in the ISAC system [53, 54]. The clusters contributed by these shared scatterers may coexist and disappear in both communication and sensing channels. Hence, the communication channel component and the sensing channel component of the ISAC channel may be partially related to each other. These novel propagation characteristics facing the ISAC technology cannot be adequately captured by existing communication channel models. Therefore, an ISAC channel model incorporating these characteristics based on realistic channel measurements is of great significance for the design and performance evaluation of ISAC systems.
I-B4 RIS, reflection coefficient characteristics, and cascaded channel
@@
RIS is considered as a potential technology that can overcome the challenges in traditional beam-space antenna array based beamforming [6]. It is capable of significantly improving the channel gain by actively controlling the wireless propagation environment [14]. For example, when the line-of-sight (LOS) path is obstructed, RIS can provide a high-power reflection path. Additionally, it promises low cost and high energy efficiency. However, since RIS can program the radio propagation environment to some degree, the role of the RIS itself should be taken into account when studying the wireless channels that involve RIS. This makes the traditional wireless channel model no longer suitable for the cascaded channel of RIS, and the channel modeling theory for RIS is not yet mature. Furthermore, for RIS, both the decoupling of the channel model and the construction of a high-precision electromagnetic response model are urgent problems to be solved [14]. In addition, the high path loss of the RIS cascaded channel and the flexibility of RIS deployment locations pose new challenges for RIS channel measurement.
@@ I-C Paper Contributions and Organization
Research on wireless channels has a vital role in the design, standardization and optimization of 6G systems. At the time of writing, a few surveys on some aspects of 6G channels [9, 16, 24, 56, 57] have been published. Nevertheless, there has not been a comprehensive survey on the entire portfolio of 6G channel research. We not only provide a comprehensive survey of the channel models of the 6G potential enabling technologies including THz, E-MIMO, ISAC, RIS and SAGIN, but also provide a tutorial on the channel simulator BUPTCMCCCMG-IMT2030, which support for various scenario and features of the above technologies. Taking into account mainstream models, the simulator uses the 5G standard channel modeling framework ITU-R M.2412 [42] (principles described in Section II) as the basis for expansion to 6G. Our major contributions via this paper are as follows:
  • The representative THz channel measurement campaigns in the 100-1100 GHz frequency band are reviewed and analyzed. Then, a detailed discussion on the THz channel characteristics is given, specifically focusing on reflection, high loss and sparsity. Finally, the pros and cons of stochastic, deterministic, and hybrid modeling approaches for THz channels are further discussed.
@@
  • The E-MIMO channel measurement platforms are systematically introduced, including frequency domain and time domain platforms. Then, the near-field spherical wave model of E-MIMO and the characteristics of SnS are discussed in detail. At the same time, the division of near-field range is analyzed. In addition, several channel modeling methods such as the statistical, deterministic and hybrid channel modeling are compared, then a realistic yet low-complexity SnS channel modeling framework is discussed. Finally, the future research directions of E-MIMO are pointed out.
  • A summary of research progress on ISAC channels in both monostatic and bistatic sensing modes is presented. Furthermore, an in-depth analysis of the novel characteristics introduced by the ISAC channel is conducted. To capture the channel sharing feature, we review a cluster-based ISAC channel model. Finally, the pros and cons of various modeling approaches are discussed for ISAC channels.
@@
  • Some measurement campaigns related to the RIS channel are summarized, especially regarding the power gain brought by RIS. In addition, the characteristics of the RIS channel are introduced in detail, including the response characteristics of the important intermediate node RIS and the new loss characteristics presented due to the cascading of two subchannels and RIS response. Then, to accurately describe the RIS channel, we provide various types of RIS channel modeling approaches, including statistical, deterministic, and hybrid approaches. Finally, future research directions are discussed.
@@
  • The 6G channel simulator BUPTCMCCCMG-IMT2030 developed by our team is introduced. The design framework and user guidelines of the simulator are detailed. We also provide simulation examples to characterize THz, E-MIMO, ISAC, RIS, and SAGIN.
@@
The rest of this paper is organized as follows. In Section III–VII, we provide an overview of the state-of-the-art channel research on THz, E-MIMO, ISAC, RIS and SAGIN technologies in terms of measurements, characteristics and modeling and discuss the prospects. Section VIII presents an tutorial on the 6G channel simulator BUPTCMCCCMG-IMT2030 developed by our team. Finally, conclusions are drawn in Section IX. For better clarity, the organization of the paper is shown in Fig. 2.
II Modeling Methodology of ITU/3GPP Stochastic Channel
The transmitted signals through the wireless channel will suffer propagation losses, which can be roughly divided into two types, large-scale and small-scale channel fading [58]. In practice, both types of fading occur [15] and can significantly affect all aspects of the communication system [16]. Large-scale fading affects the received signal strength over long Tx-Rx distances (typically hundreds or thousands of meters). Hence, it impacts coverage prediction and interference analysis, and it is an essential factor for wireless communication system deployment. However, small-scale fading generally refers to the fluctuation of signal intensity for short distances of a few wavelengths or short periods of seconds [59]. Small-scale fading is more relevant to the design of reliable and efficient communication systems [58]. The characteristics, such as the delay and the angle of MPCs, play a vital role in designing physical-layer transmission techniques [60]. In this section, we briefly introduce the modeling principles of the 5G stochastic standard channel model (ITU-R M.2412 [42] and 3GPP TR 38.901 [43]) from large-scale fading and small-scale channel fading perspectives.

### v4

# 6G Channel Modeling: Requirement, Measurement, Methodology and Simulator

@@ Abstract
Sixth-generation (6G) mobile communications have attracted substantial attention in the global research community of information and communication technologies (ICTs). 6G systems are expected to support not only extended 5G usage scenarios but also new usage scenarios, such as integrated sensing and communication (ISAC), integrated artificial intelligence (AI) and communication, and communication and ubiquitous connectivity. To achieve this goal, channel characteristics must be comprehensively studied and properly exploited to promote the design, standardization, and optimization of 6G systems. In this paper, we first summarize the requirements and challenges in 6G channel research. Our focus is on channels for six promising technologies enabling 6G, including ISAC, extremely large-scale MIMO (XL-MIMO), mid-band and terahertz (THz) technologies, reconfigurable intelligent surfaces (RISs), and space–air–ground integrated networks (SAGINs). A survey of the progress in 6G channel research regarding the above six promising technologies is presented in terms of the latest measurement campaigns, new characteristics, modeling methods, and research prospects. To support testing, optimization and evaluation, existing 6G channel simulators are summarized. Then, BUPTCMCCCMG-IMT2030 is introduced as an example of a simulator that was developed on the basis of the ITU/3GPP 3D geometry-based stochastic model (GBSM) methodology. We also address open issues covering standardization activities, AI-enabled methods, and system performance analysis in the context of 6G channel research. This paper offers in-depth, hands-on insights into the best practices of channel measurements, modeling, and simulations for the evaluation of 6G technologies, the development of 6G standards, and the implementation and optimization of 6G systems.
@@
6G, channel modeling, channel measurement, channel simulation, integrated sensing and communication, MIMO, terahertz, reconfigurable intelligent surface, space–air–ground integrated network
@@ I-A Overview of 6G
As fifth-generation (5G) mobile communication networks have been deployed worldwide, studies on sixth-generation (6G) mobile communication technologies, which are expected to be rolled out for 2030 and beyond, are being vibrantly conducted. Several 6G research projects and organizations have been established. Representative examples include the Finnish “6Genesis Flagship Program” [1], the European “Hexa-X” project [2], the Chinese “IMT-2030 (6G) Promotion Group” [3], and the North American “Next G Alliance” [4]. Simultaneously, international standardization organizations have launched several key initiatives. For example, in 2022, the International Telecommunication Union (ITU) identified future technology trends [5] for International Mobile Telecommunications (IMT) systems toward 2030 and beyond. Very recently, the recommendation framework for IMT-2030, which is commonly known as the 6G vision, was finalized [6]. This means that the global 6G standardization campaign has been formally initiated. With this recommendation framework, the ITU establishes the overall objectives, procedures and timelines for the development of 6G mobile communication standards. Following this framework, the 3rd Generation Partnership Project (3GPP) aims to complete standardization no later than 2030. As part of the technical preparation, the 3GPP has already initiated several study items in Release 20, including efforts on advanced channel modeling.
@@
6G aims to establish a more inclusive information society than 5G does. Its goal is to intelligently connect humans, machines, and various other things [6]. Some key drivers of 6G have been discussed in [5]. These drivers can be understood from the perspective associated with continuously growing mobile traffic and emerging applications. Specifically, it is forecast that by 2030, global mobile traffic will reach 5016 exabytes (EB) per month [7], which is approximately 80 times greater than that in 2020. The potential emerging applications of 6G [5] can be categorized into several main usage scenarios [6]. One extends enhanced mobile broadband (eMBB), which was introduced in 5G, to target immersive applications such as holographic communications, tactile and haptic internet applications, and extended reality (XR). Similarly, hyper-reliable and low-latency communications and massive communications are extensions of the 5G usage scenarios of ultra-reliable and low latency communications (URLLC) and massive machine type communications (mMTC), respectively. In addition, three other usage scenarios have been identified, i.e., ubiquitous connectivity, integrated artificial intelligence (AI) and communication, and integrated sensing and communication (ISAC) [6]. The applications of integrated AI/sensing and communication include perceptive mobile networks, digital twins, and the proliferation of networked AI. As a result, the capabilities of 6G are enhanced across multiple dimensions, including the data rate, spectrum efficiency, reliability, latency, mobility, and connection density [8, 9, 10, 11, 12, 13, 14, 6]. Moreover, new capabilities, such as all-domain coverage, sensing-related and AI-related capabilities, sustainability and computing power, are being considered [6].
@@ I-B Requirements and Challenges of 6G Channel Research
The radio channel is the transmission medium between a transmitter (Tx) and receiver (Rx). It plays a crucial role in determining the ultimate performance limit of mobile communications systems [15, 16, 17]. As shown in Fig. 1, 6G communication systems will incorporate various novel technological components [5, 16, 18, 10, 19, 20, 21, 8, 22], such as the integration of sensing into the communication system (i.e., ISAC) [23, 24, 25], the extension of antenna arrays to extremely large-scale multiple-input multiple output (XL-MIMO) [9, 22, 26, 27], the extension of the frequency band to new mid-band (6–24 GHz) and terahertz (THz) ranges [28, 29, 30, 31, 17, 32], the introduction of a programmable wireless environment by using reconfigurable intelligent surfaces (RISs) [33, 34, 35], and the extension of coverage to space-air-ground integrated networks (SAGINs) [36, 37, 38]. Each of these may result in distinct requirements and characteristics, as summarized below.
@@
I-B1 ISAC, target scattering, and sharing feature
@@
ISAC is one of the key technologies for 6G[39]. Compared with conventional systems with separate devices, ISAC technology can integrate functions into one system, enabling base stations (BS) or terminals to communicate while sensing the surrounding environment. In ISAC systems, sensing aims to obtain the state information, such as the speed and location, of the targets in a particular environment through echoes and the collaboration between BSs [40]. Therefore, in the study of ISAC channels, it is essential to introduce the radar cross section (RCS), a target-related feature, and consider the impact of the environment on the sensing channel [41]. Moreover, owing to the simultaneous use of the propagation environment by the communication and sensing signals, some shared propagation scatterers may be involved in the ISAC system [41, 42]. The clusters contributed by these shared scatterers may coexist and disappear in both communication and sensing channels. Hence, the communication channel component and the sensing channel component of an ISAC channel may be partially related to each other. These novel propagation characteristics facing ISAC technology cannot be adequately captured by existing communication channel models. Therefore, an ISAC channel model incorporating these characteristics on the basis of realistic channel measurements is highly important for the design and performance evaluation of ISAC systems.
I-B2 Larger-scale antenna array, near-field propagation, and spatial non-stationarity
In the 1970s, researchers suggested introducing MIMO technologies into wireless communication systems. In 1995, Telatar theoretically derived MIMO channel capacity under the fading condition [43]. It was demonstrated that the MIMO technology can substantially increase the capacity of wireless communication system. Then, the MIMO technology was introduced into the fourth generation (4G) mobile communication system, which can support eight array elements and highly improve the system spectrum efficiency [44]. However, the channel model used in 4G systems only supports the azimuth dimension. With the proposal of massive MIMO, three-dimensional (3D) MIMO [45] is a typical massive antenna realization. Therefore, the elevation dimension was introduced into the ITU/3GPP 5G standard channel model[46, 47]. The 5G channel model captures the elevation-azimuth-time-frequency channel characteristics, and meets the requirements of 3D MIMO and massive MIMO technologies. In the 6G era, the number of array elements is expected to increase significantly, reaching a range from a few hundreds to thousands. As a result, the XL-MIMO technology based on the much larger-scale antenna arrays is envisaged to be deployed. The XL-MIMO with $\times1024$ array elements can offer abundant spatial degrees of freedom for supporting 100 times peak data rate improvement in 6G mobile communication systems [48]. However, some new channel characteristics emerge in XL-MIMO systems. To elaborate a little further, when the propagation distance is shorter than the far-field Rayleigh distance, the plane wave assumption is invalid in the near-field region [26, 49]. Therefore, the the near-field spherical wave and spatial non-stationary (SnS) characteristics [31, 50] cannot be ignored in XL-MIMO communication systems. As shown in Fig. 1, the XL-MIMO can be modeled using near-field spherical propagation in the near-field region. In addition, with the expansion of physical array aperture, different array elements see different scatterers, resulting in SnS on the array. However, these characteristics are not covered by the ITU/3GPP 5G channel model framework. Therefore, it is necessary to investigate the impact of the near-field spherical wave and the SnS characteristic through channel measurements and simulations for 6G channel modeling.
I-B3 Multi-frequency, propagation attenuation, and channel sparsity
At present, the new mid-band (6–24 GHz) has attracted widespread attention and has the potential to be an important candidate frequency band for 6G. In June 2022, the 3GPP RAN Plenum adopted a standard modification proposal for the 5925–7125 MHz band, and this band was introduced into 5G-Advanced Rel-18. At the World Radiocommunication Conference 2023 (WRC-23), 6425–7125 MHz was identified for mobile use in every ITU region, and the agenda for WRC-27 was defined. The new WRC cycle will focus on the following bands for the IMT: 4400–4800 MHz, 7125–8400 MHz, and 14.8–15.35 GHz. In December 2023, the 3GPP TSG RAN Rel-19 discussed research work in the 7–24 GHz band. In [51], with the increase of frequency, channel sparsity becomes more prominent. However, the sparsity of the channel matrix derived from the 3GPP channel model is similar regardless of the frequency band, which leads to inaccurate prediction results for different frequency bands.
The deployment of 5G has already achieved highly efficient utilization of the sub-6 GHz frequency band and harnessed millimeter wave (mmWave) frequencies to deliver mobile devices with unparalleled multi-Gbps data rates [52]. The THz frequency band (0.3-3 THz according to ITU radio regulations or 0.1-10 THz as a commonly used wider range) possesses an atmospheric transmission window with low molecular absorption losses, and abundant bandwidth was allocated to terrestrial mobile communication services at the WRC in 2019 [53]. As the era of 6G begins, the demand for a peak data rate of 1 Tbps will require a comprehensive exploitation of the THz frequency band [16]. Prototype demonstration and testing systems for THz communication have already been developed to achieve high-speed transmission. For example, high-capacity wireless transmission at 1.44 Tbps have been achieved within the THz frequency band of 135-170 GHz [54]. However, it is essential to recognize that the THz band has distinctive characteristics in contrast to the sub-6 GHz and mmWave frequency bands. The shorter wavelength ($\leq$ 3 mm) signals in the THz band reflect the roughness of the objects in the environment. Previously smooth surfaces, such as walls and tree leaves, appear rougher in the THz band. The specular reflection may lose its privileged position, allowing diffuse scattering to become the predominant reflection behavior. In general, THz waves experience severe transmission losses and molecular absorption, thereby causing changes in path loss characteristics and the emergence of sparsity. Sparsity refers to the condition where only a small number of multipath components (MPCs) carry the majority of the received signal power. Currently, the ITU/3GPP channel model [46, 47] only supports frequencies up to 100 GHz for 5G. Therefore, a comprehensive THz channel model, that is capable of covering typical scenarios (such as indoor hotspot (InH) and urban microcell (UMi)) and incorporating distinctive THz channel characteristics (such as high path loss and sparsity) is urgently needed.
I-B4 RIS, reflection coefficient characteristics, and concatenated channel
@@
RIS is considered as a potential technology that can overcome the challenges in traditional beam-space antenna array based beamforming [6]. It is capable of significantly improving the channel gain by actively controlling the wireless propagation environment [14]. For example, when the line-of-sight (LOS) path is obstructed, RIS can provide a high-power reflection path. Additionally, it promises low cost and high energy efficiency. However, since RIS can program the wireless propagation environment to some degree, the role of the RIS itself should be taken into account when studying the wireless channels that involve RIS. This makes the traditional wireless channel model no longer suitable for the concatenated channel of RIS, and the channel modeling theory for RIS is not yet mature. Furthermore, for RIS, both the decoupling of the channel model and the construction of a high-precision electromagnetic response model are urgent problems to be solved [14]. In addition, the high path loss of the RIS concatenated channel and the flexibility of RIS deployment locations pose new challenges for RIS channel measurement.
@@ I-C Paper Contributions and Organization
Research on wireless channels plays a vital role in the design, standardization and optimization of 6G systems. At the time of this writing, a few surveys on some aspects of 6G channels [9, 16, 29, 56, 57] have been published. Nevertheless, there has not been a comprehensive survey of the entire portfolio of 6G channel research. We provide a comprehensive survey of the channel models of potential 6G enabling technologies, including ISAC, XL-MIMO, mid-band and THz technologies, RISs and SAGINs. 6G channel simulators supporting potential enabling technologies are summarized, and the Beijing University of Posts and Telecommunication and China Mobile Communications Corporation Channel Model Generator-IMT2030 (BUPTCMCCCMG-IMT2030) is introduced as an example, which supports various features of potential 6G technologies. Considering mainstream models, the simulator uses the 5G standard channel modeling framework ITU-R M.2412 [46] (principles described in Section II) as the basis for expansion to 6G. Our major contributions via this paper are as follows:
  • An overview of ISAC channel measurements is provided, including measurement platforms based on time-domain correlation and vector network analyzers (VNAs), as well as monostatic and bistatic sensing modes. Then, an in-depth analysis of the new channel characteristics introduced by ISAC is conducted, with a particular focus on target scattering, EO, Doppler shift, background effect, and sharing feature. A cluster-based ISAC channel model is introduced to characterize these characteristics accurately. Finally, the pros and cons of various ISAC channel modeling methods are discussed.
@@
  • The XL-MIMO channel measurement platforms, including frequency domain and time domain platforms, are systematically introduced. Next, a near-field spherical wave model of XL-MIMO and the characteristics of SnS are discussed in detail. At the same time, the division of near-field range is analyzed. In addition, several channel modeling methods such as the statistical, deterministic and hybrid channel modeling are compared, then a realistic yet low-complexity SnS channel modeling framework is discussed. Finally, the future research directions of XL-MIMO are pointed out.
  • The representative multi-frequency channel measurement campaigns, covering both mid-band and THz frequencies, are reviewed and analyzed. Then, a detailed discussion on the multi-frequency channel characteristics is given, specifically focusing on reflection, high loss and sparsity. Finally, the pros and cons of stochastic, deterministic, and hybrid modeling approaches for multi-frequency channels are further discussed.
@@
  • Some measurement campaigns related to the RIS channel are summarized, especially regarding the power gain brought by RIS. In addition, the characteristics of the RIS channel are introduced in detail, including the response characteristics of the important intermediate node RIS and the new loss characteristics presented due to the concatenation of two subchannels and RIS response. Then, to accurately describe the RIS channel, we provide various types of RIS channel modeling approaches, including statistical, deterministic, and hybrid approaches. Finally, future research directions are discussed.
@@
  • 6G channel simulators that support potential enabling technologies are summarized, and the BUPTCMCCCMG-IMT2030 simulator developed by our team is introduced, including its principles, structure, functions, and applications.
@@
The rest of this paper is organized as follows. In Section IIIVII, we provide a survey of the state-of-the-art channel research on ISAC, XL-MIMO, mid-band and THz technologies (as a multi-frequency study), RISs and SAGINs in terms of measurements, characteristics and modeling and discuss the prospects. Section VIII summarizes 6G channel simulators that support potential enabling technologies and introduces BUPTCMCCCMG-IMT2030 as a standards-based implementation. Section IX presents open issues in 6G channel research, including recent 3GPP studies on channel modeling for 7–24 GHz bands and ISAC, as well as AI-enabled methods in channel research and system performance analysis for practical channel models. Finally, conclusions are drawn in Section X. For clarity, the organization of the paper is shown in Fig. 2.
II Modeling Methodology of the ITU/3GPP Stochastic Channel
This section introduces the standardized 5G channel models defined in 3GPP TR 38.901 [47] and ITU-R M.2412[46]. These models serve as the foundational framework upon which our proposed 6G channel extensions are built. Specifically, the new characteristics and modeling approaches discussed in Sections III to VII—including those for ISAC, XL-MIMO, and other key 6G technologies—are all based on extensions to this baseline model.
In wireless communications, transmitted signals are subject to propagation losses, which are broadly classified into large-scale and small-scale fading effects [58]. In practice, both types of fading coexist [15] and can significantly impact various aspects of communication system design and performance [16]. Large-scale fading affects signal strength over long Tx–Rx distances (typically hundreds or thousands of meters,) and is crucial for coverage prediction and interference analysis, thus playing a vital role in wireless system deployment. On the other hand, small-scale fading refers to rapid signal fluctuations over short distances (on the order of a few wavelengths) or brief time intervals [59]. This type of fading is particularly important for the design of reliable and efficient physical-layer transmission techniques [58]. The delay and angular characteristics of multipath components (MPCs), for instance, are critical for optimizing channel estimation, beamforming, and MIMO performance [60]. This section outlines the modeling principles of the 5G stochastic channel model from both large-scale and small-scale perspectives, forming the basis for subsequent 6G model development.